6. On a body, which has a axis of rotation, act two pairs of forces, one right with value M1 = + 50 Nm and the other left with value M2 = - 20 Nm. a) Calculate the resultant moment. b) Can this body be in balance? Why ? c) If the body is not in equilibrium, in what direction will it rotate, in that of the right or left moment?


1
Expert's answer
2021-03-09T15:30:35-0500

(a) Let's calculate the resultant moment:


"M_{net}=M_1+M_2=50\ N\cdot m+(-20\ N\cdot m)=30\ N\cdot m."

(b) If the body is in balance the momenta must be equal to each other. But, the body isn't in equilibrium because "M_1 \ne M_2."

(c) The resultant moment has the sign plus which means that the body will rotate in the direction of the right moment.